OpenStudy (anonymous):

The student government can sell a total of 400 tickets to a basketball game. An adult ticket costs $3.00 and a student ticket costs $2.00. What is the minimum number of adult tickets that must be sold if the student government wants to collect at least $1100.00? Thanks for any help!

OpenStudy (anonymous):

300 adult tickets

OpenStudy (tylerd):

1100 = 3a +2s a+s=400

OpenStudy (texaschic101):

a + s = 400 3a + 2s > 1100

OpenStudy (texaschic101):

oops.. 3a + 2s >= 1100